OpenStudy (anonymous):

Which is the 25th term of the arithmetic sequence where a1 = 7 and a9 = 47

Parth (parthkohli):

The common difference is 40/8 = 5 The formula: \(\Large \color{Black}{\Rightarrow a_{n} = a_{1} + (n - 1)d }\) Here: \(\Large \color{Black}{\Rightarrow a_{25} = 7 + 5(24 - 1) }\)

OpenStudy (cwrw238):

first find common difference = ( a9 - a1) / 8 then use formula for 25the term

OpenStudy (anonymous):

is it 122??

Parth (parthkohli):

Yes.

Parth (parthkohli):

:P

OpenStudy (anonymous):

but thats not in my answers:(

OpenStudy (cwrw238):

7 + 5*24

Parth (parthkohli):

Wouldn't it be 5 * 23?

OpenStudy (cwrw238):

25th term = a1 + 24d

OpenStudy (anonymous):

its 127,

Parth (parthkohli):

Oh sorry lol

OpenStudy (cwrw238):

right

Parth (parthkohli):

(n - 1)d

Parth (parthkohli):

24 * d

OpenStudy (cwrw238):

yea

Parth (parthkohli):

Lol I got confused by myself

OpenStudy (cwrw238):

that happens - none of us are computers!!!

Parth (parthkohli):

Hehe lol thanks

OpenStudy (anonymous):

haha thanks can you help me with this ,What is the 8th term of the geometric sequence where a1 = 1024 and a3 = 64, and its no worries (:

OpenStudy (cwrw238):

i can but the moderators like you to post another question separately

OpenStudy (cwrw238):

best to do that, i think